

********************************************************************
UARS WEEKLY STATUS REPORT - 03 AUG 2001
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, taking solar data on all available
orbital tracking periods and stellar calibration observations whenever
possible. We continue to gather enough data to meet all the SOLSTICE
science objectives during this period of reduced tape recorder operations.
***********
SUSIM
***********
Thanks a generous donation of power by PEM, we were able to plan for
nearly all of our final (?) lamp calibration. We are most grateful.
The sequence of corresponding reference channel solar scans begins today.
SUSIM has completed its daily solar scans this past week during
periods of predicted UARS telemetry.
***********
ACRIM
***********
Nothing received.
***********
PEM
***********
Nothing received.
***********
HALOE
***********
Nothing received.
***********
MLS
***********
Nothing received.
***********
HRDI
***********
Nothing received.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
MPG distributed a revised copy of the Instrument Schedule
for the next yaw period. On 01 August the MPG participated
in the UARS Monthly Status Review meeting.
Daily processing of instrument Daily Activity Plans
continues with no problems.
***********
CDHF
***********
Nothing received.
***********
FOT
***********
REPORTING PERIOD
This report covers 21 July 2001 (Orbit 53896, GMT Day 202) through 27 July 2001
(Orbit 54000, GMT Day 208).
SPACECRAFT OPERATION
The observatory is now in Forward Flight (Southern Hemisphere viewing) and is
performing nominally under two-battery operations. The instrument operational
changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
PEM 202/0746 53900 Power Off
PEM 202/0929 53901 Day Only
PEM 202/1526 53905 Full Time
PEM 202/1849 53907 Day Only
PEM 202/2029 53908 Power Off
PEM 202/2350 53910 Day Only
HALOE 208/0004 53986 Power On
PEM 208/0610 53898 AXIS On
The beta angle ranged between the angles of 75.7 degrees to 54.1 degrees for
this report period. The beta angle is now decreasing towards a minimum angle
of 0 degrees on 09 August 2001 (DOY 221).
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 23.8 V end-of-night (EON) load bus voltage. Battery
temperatures are stable with a temperature delta between Battery 2 and Battery
3 of 1.2 to 1.3 degrees C. Additional battery performance data for each
battery on 20 July 2001 and 27 July 2001 are provided in Appendix D.
The clock error ranged from Ð11.3 to 6.3 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial excursion of
ALERTS or Out-Of-Limits conditions detected by the FOT for this report. A
summary is listed below and the details of ALERT and Out-Of-Limits occurrences
are listed in Appendix B.
¥ OUT-OF-LIMITS = 02
¥ ALERTS = 00
The following real-time command sheets were executed for the listed operational
element(s) during this report period (see Appendix A).
¥ FOT = 04
¥ HALOE = 01
¥ PEM = 09
UARS Anomaly Reports initiated during this report period are listed in Appendix
E and summarized below.
¥ FOT = 01
One (1) Anomaly Report remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
· Transponder A: 00/316 - 2287.496493
· Transponder B: 01/119 - 2287.499116
GROUND SYSTEM OPERATION
One (1) operational support problems occurred this period. Anomaly Reports are
included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from mission
start through UTC Day 01/207 show the observatory data loss to be 302 hours, 55
minutes, 48.256 seconds. This shows a loss of 06 minutes and 13.088 seconds
since the last report period. This is a 0.39 percent data loss, which equals a
99.61 percent data capture for the mission.
The increase of 06 minutes and 13.088 seconds occurred over five (5) days:
· 35.168 seconds on Day 202 due to a Data Hit.
· 1 minute and 16.256 seconds on Day 204 due to Data Hits
· 4 minutes and 08.000 seconds on Day 205 due to bad weather TTR# 23977
· 13.664 seconds on Day 207 due to a Data Hit
OTHER SIGNIFICANT EVENTS
UARS SATELLITE OPERATIONS
The UARS FOT attended a UARS All-Hands Meeting on Wednesday, July 25,2001. At
this meeting the FOT and all UARS entities were informed that NASA headquarters
has stated (verbally) that there will be no funds available to operate UARS
past this September. This statement is unofficial at this time, but the FOT
has begun work on the decommissioning plan, which will end all UARS operations.
All development activities will be continued until this decision is finalized;
however, they may be discontinued pending the outcome from NASA headquarters.
The FOT changed the safehold configuration on board this week. The previous
configuration had been sun-pointing safehold, which only gave a few orbits of
time before the batteries would suffer severe overcharging and most likely end
the mission. The current configuration is now earth-pointing safehold with the
earth sensor off. This will allow the spacecraft to keep pitching over at one
revolution per orbit and keep the other axis stable for a minimum of 3-4
orbits. Given the almost realtime status of UARS operations this should give
enough time to react to a safehold situation.
MEETINGS
The UARS FOT attended a UARS All-Hands Meeting on Wednesday, July 25,2001.
The UARS FOT attended a UARS SERB/TAC RS Meeting on Wednesday, July 25,2001.
The UARS FOT attended a Safehold Working Group Meeting on Thursday, July 26,
2001.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw slew is scheduled for 08 August 2001 (DOY 220). This
maneuver will be a Forward to Reverse Flight maneuver (Southern Hemisphere
viewing).
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 27 July 2001
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ ------
53900 202/0746 01-465 PEM (Z/H) Power OFF
53901 202/0929 01-466 PEM (Z/H) Power ON DAY ONLY
53905 202/1526 01-467 PEM (Z/H) Power ON FULL TIME
53907 202/1849 01-468 PEM (Z/H) Power ON DAY ONLY
53908 202/2029 01-469 PEM (Z/H) Power OFF
53910 202/2350 01-470 PEM (Z/H) Power ON DAY ONLY
53911 203/0106 01-471 PEM (Z/H) TO FULL TIME OPS
53924 203/2132 01-472 OBC/CLK Clk Adj from 4.20 to 20.6 msec/day
53980 207/1449 01-473 MACS Safehold Reconfig
53982 207/1809 01-474 NBTR NOOPS D-208
53986 208/0004 01-475 PEM ZEPS HEPS TO DAY ONLY
53986 208/0004 01-476 HALOE Power ON
53898 208/0610 01-477 PEM Power ON AXIS
53997 208/1810 01-478 NBTR D209 noops
______________________________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 27 July 2001
The following Out-Of-Limits occurrences were detected:
MNEMONIC STATE DESCRIPTION ORBITS
-------- ----- ----------- ------
CDUNRG28 Y-LO CDH UNREGULATED 28V 53900
HRPSEQBIMON Y-HI HRDI QB CURRENT MONITOR 53899,900
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
-------- ----------- ------
-None this report period
_______________________________________________________________________________
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 27 July 2001
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ----- -----------------
53924 203/2132 Clk Adj from 4.20 to 20.6 msec/day
_______________________________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 27 July 2001
20 July 2001 - GMT Day 201 Beta = 77.4 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- -------- ---------- ----- ------- ------
1 +728.0/+728.0 -1.79/-3.05 +0.0/+0.0 12.2 0.000 0.0
2 +11.2/-89.6 +2.60/+1.34 +14.4/-15.2 31.5 0.096 13.3
3 +16.8/-44.8 +1.97/+1.03 +14.0/-15.6 31.5 0.095 13.3
27 July 2001 - GMT Day 208 Beta = 54.1 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------- ---------- ---------- ----- ------- ------
1 +728.0/+728.0 -3.68/-4.94 +0.0/+0.0 12.0 0.000 0.0
2 +11.2/-67.2 +4.51/+3.23 +24.8/-15.2 25.0 1.050 17.9
3 +22.4/-44.8 +3.55/+2.29 +24.0/-14.4 24.8 1.050 17.0
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- ---- ---------
-no changes V/T 5
______________________________________________________________________________
APPENDIX E
UARS ANOMALY REPORTS
Week ending 27 July 2001
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-044 FOT 53941 205/0004 23997 Late Acq. due to
weather at W.S.
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-044 FOT 53941 205/0004 23997 Late Acq. due to
weather at W.S.
ANOMALY INVESTIGATION REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
-None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
01-041 FOT see list July 01 14170 Generic Late
Acquisition
SEVNTFW01/30